Jefferson Lab HALL A Logbook for May 00 - All EXCEPT Automatic Entries

  • Main INDEX

    Wednesday

  • 42592 05/31/00 22:32 Olivier straw efficiency (rear chambers)
  • 42590 05/31/00 22:30 Olivier straw efficency (front chambers)
  • 42570 05/31/00 20:02 M Jones access to fix fpp
  • 42569 05/31/00 19:34 Olivier fpp demux pb
  • 42554 05/31/00 18:17 Olivier harp scan
  • 42545 05/31/00 17:14 R. Michaels re: double self timing peaks
  • 42541 05/31/00 16:55 Calarco shift summary
  • 42507 05/31/00 13:03 R. Michaels trigger tests
  • 42485 05/31/00 11:29 R. Michaels self timing peak doubling
  • 42484 05/31/00 11:24 Calarco H-arm TDC problem
  • 42477 05/31/00 10:58 Calarco S2 Hadron TDCs
  • 42460 05/31/00 08:04 R Roche electron arm HV
  • 42458 05/31/00 08:02 R Roche shift summary
  • 42435 05/31/00 02:05 R Roche FPP straw efficiencies - part II
  • 42434 05/31/00 02:01 R Roche FPP straw efficiencies
  • Tuesday

  • 42421 05/30/00 23:58 Breuer shift summary
  • 42381 05/30/00 18:50 Punjabi harp scan
  • 42368 05/30/00 15:56 Kumbartzki day shift summary
  • 42367 05/30/00 11:29 milbrath straw fpp
  • 42350 05/30/00 08:03 Kelly owl shift summary
  • 42349 05/30/00 07:20 Kelly missing straws in fppFV1,2,3
  • 42332 05/30/00 04:53 Kelly magnet screen at end of thcm=-90
  • 42313 05/30/00 01:58 Kelly rear straw efficiency
  • 42312 05/30/00 01:56 Kelly front straw efficiency
  • 42308 05/30/00 00:37 breuer shift summary
  • Monday

  • 42299 05/29/00 21:47 Breuer target reboot manual heater
  • 42284 05/29/00 16:29 h . breuer harp scan result
  • 42279 05/29/00 16:00 kumbartzki day shift summary
  • 42262 05/29/00 09:26 kumbartzki UPS software crash
  • 42255 05/29/00 08:01 Kelly owl shift summary
  • 42254 05/29/00 07:22 Kelly magnet screen for thcm=-90
  • 42236 05/29/00 04:38 Kelly magnet screen, thcm = -50
  • 42231 05/29/00 04:13 Kelly harp scan 915
  • 42220 05/29/00 00:02 M. Rvachev Shift summary
  • Sunday

  • 42211 05/28/00 19:46 Breuer target IOC reboot
  • 42202 05/28/00 15:46 G. Kumbartzk day shift summary
  • 42195 05/28/00 14:09 kumbartzki harp scan
  • 42193 05/28/00 13:30 kumbartzki fpp efficiency back chambers
  • 42192 05/28/00 13:24 kumbartzki fpp efficieny front
  • 42186 05/28/00 09:32 Kelly elastic tail in thcm=50,phi=180
  • 42179 05/28/00 08:02 Kelly owl shift summary
  • 42159 05/28/00 06:17 Kelly harp scan plot, 911
  • 42130 05/28/00 00:04 milbrath shift summary
  • Saturday

  • 42106 05/27/00 20:24 Meek Beam down for reorbit
  • 42105 05/27/00 20:23 Adaq HARP scan beam pos 0,0
  • 42069 05/27/00 16:32 Adaq harp
  • 42064 05/27/00 16:10 G.Kumbartzki shift sumary
  • 42062 05/27/00 16:06 Punjabi harp scan
  • 42056 05/27/00 12:15 meek New HRSe rotation limit.
  • 42055 05/27/00 12:13 meek target cool down
  • 42054 05/27/00 11:10 M Jones LVDT settings after target work
  • 42053 05/27/00 10:19 Meek Cool down started
  • 42052 05/27/00 10:01 Meek Cryotarget Cool Down start
  • Friday

  • 42050 05/26/00 10:49 mitchell LVDT's before target work
  • 42049 05/26/00 09:08 Kumbartzki Day shift summary
  • 42047 05/26/00 08:03 M JOnes shift summary owl shift
  • 42045 05/26/00 05:47 M JOnes IN access
  • 42040 05/26/00 03:03 M Jones possible energy fluctuations
  • 42039 05/26/00 03:01 M Jones harp scan
  • 42028 05/26/00 00:30 Perdrisat log for 16-24 shift
  • Thursday

  • 41998 05/25/00 15:58 mitchell day shift summary
  • 41963 05/25/00 11:21 mitchell trigger stability T5's
  • 41962 05/25/00 11:11 mitchell trigger stability-T1s
  • 41953 05/25/00 07:58 M Jones owl shift summary
  • 41935 05/25/00 01:57 Meek Normalized T5 triggers
  • 41930 05/25/00 00:02 LeRose Shift Summary
  • Wednesday

  • 41912 05/24/00 16:10 B. Beise Shift summary
  • 41906 05/24/00 13:56 mitchell T1 stability
  • 41905 05/24/00 13:54 mitchell trigger stability
  • 41877 05/24/00 08:09 M. Rvachev "ref histograms" feature of dhist
  • 41875 05/24/00 08:00 R Roche shift summary
  • 41872 05/24/00 07:41 M. Rvachev Shower block defective?
  • 41848 05/24/00 00:06 MILBRATH shift
  • Tuesday

  • 41828 05/23/00 20:41 Eugene Moller results
  • 41827 05/23/00 17:21 Higinbotham Arc Energy Measurement
  • 41825 05/23/00 15:44 Mitchell shift summary
  • 41809 05/23/00 13:06 Adaq target luminosity scan -- scaler analysis
  • 41804 05/23/00 08:11 R Roche shift summary
  • 41783 05/23/00 00:07 sarty shift summary
  • Monday

  • 41780 05/22/00 22:58 sarty unlack of BPM data
  • 41779 05/22/00 22:43 Kelly lack of BPM data
  • 41773 05/22/00 21:29 Ed Brash ssh login to adaq accounts
  • 41771 05/22/00 21:23 M Jones fpp fixed
  • 41769 05/22/00 20:05 M JOnes access to fix fpp problem
  • 41743 05/22/00 16:21 Saha Harp scanners all working
  • 41742 05/22/00 16:02 bogdanw shift summary
  • 41740 05/22/00 15:59 sarty magnet screen
  • 41737 05/22/00 14:31 mitchell ep energy measurement
  • 41716 05/22/00 11:03 jpchen beam back
  • 41714 05/22/00 09:54 markowitz beam gone
  • 41710 05/22/00 09:43 markowitz fpp hv screen
  • 41709 05/22/00 09:43 markowitz magnets
  • 41687 05/22/00 07:59 R. Roche Shift summary
  • 41674 05/22/00 06:23 Adaq front FPP eff for run 1175
  • 41653 05/22/00 03:57 M Jones Time estimate for Q2=1.37 Ac measurement
  • 41652 05/22/00 03:48 M Jones Rear chamber straw efficiencies run 1161
  • 41651 05/22/00 03:47 M Jones run 1161 straw efficiency front chambers
  • Sunday

  • 41627 05/21/00 23:55 sarty Shift Summary
  • 41624 05/21/00 22:33 sarty BCM Calibration run
  • 41610 05/21/00 20:39 Adaq magnet Q2=4.13
  • 41607 05/21/00 19:06 milbrath gas chart
  • 41605 05/21/00 18:44 sarty raster ioc reboot
  • 41598 05/21/00 17:45 sarty magnet screen for H elastic (Q^2=3.48)
  • 41592 05/21/00 16:24 Adaq magnet screen elastic Q^2=2.65
  • 41585 05/21/00 16:01 markowitz shift summary
  • 41580 05/21/00 15:15 Saha Harp Scans - Raster Size
  • 41566 05/21/00 13:05 Adaq halog problems/jlabs2 disk quota
  • 41563 05/21/00 12:11 saw Hall A log book problems
  • 41562 05/21/00 12:06 Punjabi harp scan
  • 41559 05/21/00 11:58 Punjabi harpp scan
  • 41554 05/21/00 11:42 ??? A picture
  • 41541 05/21/00 10:30 ??? A picture
  • 41540 05/21/00 10:23 ??? Lost entry
  • 41514 05/21/00 08:04 M Jones Shift Summary
  • 41479 05/21/00 01:46 Meek Harp swipe for luminosity scan
  • Saturday

  • 41460 05/20/00 23:59 sarty Shift Summary
  • 41454 05/20/00 23:25 Adaq magnet screen target luminosity scans
  • 41440 05/20/00 20:58 Adaq magnet screen -2%delta
  • 41418 05/20/00 16:42 Adaq magnet screen -2% delta
  • 41413 05/20/00 16:10 Kelly shift summary
  • 41391 05/20/00 13:40 sarty re-measure HARP/BPM scans
  • 41349 05/20/00 09:37 M Jones Electron Q2 wrong polarity runs 1025-1032
  • 41342 05/20/00 07:49 adaq beam trip
  • 41341 05/20/00 07:38 adaq run 1032 is LH2
  • 41333 05/20/00 06:12 adaq sieve out
  • 41325 05/20/00 05:02 adaq run 1025
  • Friday

  • 41322 05/19/00 18:55 milbrath Q3 trip
  • 41320 05/19/00 18:17 Milbrath Q3 trip
  • 41318 05/19/00 17:35 meek Summary of Raster scan:
  • 41317 05/19/00 17:23 M Jones Shift summary
  • 41316 05/19/00 17:09 meek Requested access to fix Q3
  • 41315 05/19/00 17:06 meek harp scan raster 1x1 file 873
  • 41311 05/19/00 16:59 meek Harp scan raster 2x2 file 872
  • 41308 05/19/00 16:38 meek Harp scan file 867 pos:(0.026,3.610)
  • 41304 05/19/00 16:04 meek Harp scan file 866 pos:(-0.021,-5.123)
  • 41303 05/19/00 15:53 meek Harp scan file 865 pos (-4.493,-0.147)
  • 41302 05/19/00 15:47 meek Harp scan file 864 pos: (4.471,-0.402)
  • 41301 05/19/00 15:36 meek Harp scan file 863 pos:(0.0,-0.3)
  • 41300 05/19/00 15:28 meek Harp scan file number 862
  • 41294 05/19/00 08:35 Kelly shift summary
  • Thursday

  • 41291 05/18/00 21:31 meek Target cooldown and survey.
  • Wednesday

    
    

    Tuesday

  • 41270 05/16/00 15:09 roblin Ep result final for E=0.885
  • 41260 05/16/00 09:33 M Jones fpp chamber 1 fix
  • Monday

  • 41258 05/15/00 09:25 M Jones fpp gas
  • Saturday

  • 41257 05/13/00 11:21 kuss target is warming up
  • 41247 05/13/00 08:00 adaq delta=4% settings
  • 41226 05/13/00 01:33 adaq delta=0% settings
  • 41205 05/13/00 00:05 Kelly shift summary
  • Friday

  • 41198 05/12/00 23:45 Adaq delta=+2% magnet screen
  • 41195 05/12/00 23:36 Kelly reluctant p0 responses
  • 41163 05/12/00 15:55 S. Frullani Shift Summary
  • 41138 05/12/00 13:05 Eugene Moller results
  • 41124 05/12/00 08:03 M Jones owl shift summary
  • 41082 05/12/00 04:30 M Jones spot ++
  • 41079 05/12/00 04:16 B. Hu Fpp gas
  • 41076 05/12/00 04:13 B. Hu magnet screen
  • 41070 05/12/00 01:15 R. Michaels rate vs current - lead test
  • 41069 05/12/00 01:14 R. Michaels current ramp-up continued
  • 41066 05/12/00 01:03 R. Michaels trigger/current for lead vs time
  • Thursday

  • 41038 05/11/00 23:14 R. Michaels elastic at x = -2.0
  • 41037 05/11/00 22:38 R. Michaels ramp up of current on lead
  • 41030 05/11/00 22:18 R. Michaels momentum at x=2 mm
  • 41028 05/11/00 22:14 R. Michaels momentum spectrum
  • 41022 05/11/00 22:00 R. Michaels put sieve slit in (E-arm)
  • 41016 05/11/00 21:43 R. Michaels lead tests -- beam scan
  • 41003 05/11/00 20:59 R. Michaels magnets/epics screen - lead
  • 41001 05/11/00 20:56 R. Michaels check beam with BeO (useless !)
  • 40996 05/11/00 20:47 R. Michaels spot size for lead tests
  • 40986 05/11/00 20:29 R. Michaels restore trigger, move spectrom.
  • 40966 05/11/00 18:02 M Jones Restored HV epics readout
  • 40962 05/11/00 16:01 Olivier end of LG test
  • 40951 05/11/00 10:03 Olivier FFB
  • 40946 05/11/00 08:18 M Jones owl shift summary
  • 40924 05/11/00 00:01 gilman evening shift summary
  • Wednesday

  • 40904 05/10/00 16:02 S. Strauch Day shift summary
  • 40902 05/10/00 15:29 R. Michaels SPOT calibration mod
  • 40884 05/10/00 08:16 M JOnes owl shift summary
  • 40878 05/10/00 06:33 M Jones change to detmap_lg.config
  • 40873 05/10/00 05:58 M Jones scan of adc gate
  • 40865 05/10/00 04:49 M Jones access to halla
  • 40864 05/10/00 04:47 B. HU harp again
  • 40855 05/10/00 03:45 B. HU harp
  • 40854 05/10/00 03:40 B. HU OTR is removed
  • 40850 05/10/00 03:09 B. HU electron adc gate change and pedrun
  • 40841 05/10/00 01:22 B. HU electron scintinator OFF
  • 40840 05/10/00 01:19 Eugene Moller results
  • Tuesday

  • 40837 05/09/00 23:58 punjabi, gil evening shift summary
  • 40831 05/09/00 21:10 Olivier move target
  • 40830 05/09/00 20:59 Olivier prescalers
  • 40828 05/09/00 15:07 jpchen Cryotarget switched to LH2
  • 40827 05/09/00 14:32 Olivier Set hadron spectrometer
  • 40826 05/09/00 12:06 R. Michaels ET/Fortran performance
  • 40800 05/09/00 10:33 R. Michaels DAQ tests
  • 40799 05/09/00 10:33 R. Michaels trigger mods, Pedrisat's test
  • Monday

  • 40769 05/08/00 23:48 sarty Shift summary
  • 40758 05/08/00 16:11 G. M. Urciuo Day shift
  • 40747 05/08/00 13:13 krishniw magnet screens
  • 40744 05/08/00 12:27 kuss 4He heating studies
  • 40713 05/08/00 09:21 Ulmer Summary of injector energy measurements
  • 40711 05/08/00 08:05 tony Owl Shift Summary
  • 40703 05/08/00 04:36 tony ROC 14 reset (03:30)
  • 40702 05/08/00 04:23 tony TS1 rebooted (02:15)
  • 40694 05/08/00 00:06 sarty Shift summary
  • Sunday

  • 40693 05/07/00 23:58 sarty Helicity readout: tiny glitch?
  • 40692 05/07/00 23:54 Kelly magnet screen
  • 40691 05/07/00 23:53 Kelly fpp gas charts
  • 40687 05/07/00 17:22 sarty BPM Pedestal Run
  • 40678 05/07/00 16:09 G. M. Urciuo Day shift
  • 40677 05/07/00 15:16 Kelly magnet screen
  • 40675 05/07/00 15:15 Kelly fpp gas screen
  • 40672 05/07/00 11:10 Ulmer Statistics estimate
  • 40670 05/07/00 09:44 Kelly T5 total, runs 1951-2093
  • 40667 05/07/00 08:05 tony Owl Shift Summary
  • 40665 05/07/00 07:43 tony Beam is back
  • 40663 05/07/00 06:53 tony S. linac crash
  • 40661 05/07/00 06:02 tony RF Trip S. linac
  • 40654 05/07/00 03:50 tony FPP gas screen
  • 40651 05/07/00 01:46 tony Magnet Screen
  • 40650 05/07/00 01:45 tony Beam Returns
  • 40649 05/07/00 01:40 Mitzi Spot Check
  • 40647 05/07/00 01:02 tony Beam Status
  • 40645 05/07/00 00:08 Higinbotham Shift Summary
  • Saturday

  • 40643 05/06/00 23:31 Higinbotham Refresh Q1 Screen
  • 40640 05/06/00 22:55 Higinbotham FPP Gas Strip Chart
  • 40639 05/06/00 22:35 Higinbotham BPM medm Screens
  • 40638 05/06/00 22:30 Higinbotham BPM Pedestal Run
  • 40633 05/06/00 21:09 Higinbotham 10cm Dummy Cell
  • 40628 05/06/00 19:13 Higinbotham Target IOC Problem
  • 40627 05/06/00 19:10 Higinbotham Carbon Pointing
  • 40625 05/06/00 18:59 y. r. roblin EP measurement
  • 40617 05/06/00 15:59 A. Gasparian beam
  • 40603 05/06/00 08:18 tony Owl Shift Summary
  • 40596 05/06/00 04:46 tony Beam Returns
  • 40595 05/06/00 04:42 tony spot check after down time
  • 40593 05/06/00 04:14 tony beam update
  • 40592 05/06/00 01:52 tony Beam Down
  • 40590 05/06/00 00:27 tony fpp gas stripchart
  • 40589 05/06/00 00:25 tony Owl Shift Magnet Screen
  • 40585 05/06/00 00:04 Higtinbotham Shift Summary
  • Friday

  • 40580 05/05/00 20:53 Higinbotham Arc Energy Measurement
  • 40579 05/05/00 20:43 Higinbotham FPP Gas striptool log
  • 40575 05/05/00 18:18 Eugene Moller results
  • 40574 05/05/00 16:01 gilman day shift summary
  • 40562 05/05/00 12:24 kuss bulleye scan
  • 40561 05/05/00 12:24 Saha BPM/Harp Calibration
  • 40532 05/05/00 10:59 Ulmer Injector energy measurement
  • 40530 05/05/00 10:20 Ulmer Injector energy measurement bogus
  • 40524 05/05/00 08:09 urciuoli owl shift summary
  • 40517 05/05/00 02:36 Gueye How to move the Compton detector out
  • 40511 05/05/00 00:02 Higinbotham Shift Summary
  • Thursday

  • 40498 05/04/00 20:35 Higinbotham fpp gas
  • 40497 05/04/00 19:41 Higinbotham Spot++
  • 40489 05/04/00 18:04 R. Michaels more accurately, run 2027
  • 40488 05/04/00 17:26 R. Michaels BPM study, spot calibration
  • 40487 05/04/00 16:15 ulmer Beam position study
  • 40485 05/04/00 16:03 ransome shift summary
  • 40480 05/04/00 15:29 Ulmer Injector energy measurement
  • 40476 05/04/00 13:02 sabatie ROC 1 2 3 14 crashed
  • 40471 05/04/00 11:25 ransome spot change
  • 40470 05/04/00 11:14 Ransome cryogenic problems
  • 40459 05/04/00 04:35 Gueye Cavity problem
  • 40454 05/04/00 01:27 Gueye OTR out
  • 40453 05/04/00 01:18 Gueye OTR test during run 2015
  • Wednesday

  • 40447 05/03/00 23:10 sabatie Compton Polarimeter scan
  • 40446 05/03/00 22:17 ransome roc problems, runs 2011,2012
  • 40443 05/03/00 21:52 Adaq fpp strip chart
  • 40441 05/03/00 20:19 Eugene Moller results
  • 40431 05/03/00 15:21 ulmer BPM3B software offset?
  • 40430 05/03/00 15:12 ulmer Estimate of beam restoration time
  • 40427 05/03/00 05:42 Olivier Reset Q1 display
  • 40426 05/03/00 03:19 Olivier Reset Q1 display
  • Tuesday

  • 40424 05/02/00 17:24 M Jones fpp fix
  • 40422 05/02/00 15:29 jpchen controlled access
  • 40419 05/02/00 09:14 jpchen Target moved out of beam
  • 40417 05/02/00 07:02 Gueye Q1H back.
  • 40415 05/02/00 06:53 Gueye Q1H tripped
  • 40413 05/02/00 05:55 Gueye Notes on run2004 Ctnd ...
  • 40411 05/02/00 05:47 Gueye Notes on run2004
  • 40409 05/02/00 05:44 Gueye MCC OTR test
  • 40402 05/02/00 01:08 gueye Compton test
  • Monday

  • 40397 05/01/00 23:43 Hyde-Wright Swing Shift Summary
  • 40396 05/01/00 23:37 Hyde-Wright compton
  • 40392 05/01/00 23:06 Sabatie Target
  • 40385 05/01/00 19:10 Hyde-Wright fpp gas
  • 40382 05/01/00 18:53 Tony Latest Straw Eff.
  • 40375 05/01/00 15:35 Olivier back to data
  • 40372 05/01/00 15:16 Olivier No acquisition
  • 40366 05/01/00 14:29 Olivier e-D1 and h-Q1 trips
  • 40364 05/01/00 13:53 Olivier electrole dipole tripped
  • 40361 05/01/00 12:22 J. Nappa reset Q1 display
  • 40360 05/01/00 12:08 Olivier beam back
  • 40359 05/01/00 12:02 Olivier beam stop
  • 40356 05/01/00 10:45 Olivier beam back
  • 40354 05/01/00 10:19 Olivier injector in controlled access
  • 40352 05/01/00 09:45 Olivier recesiation
  • 40351 05/01/00 08:45 Olivier more current
  • 40350 05/01/00 08:29 Olivier beam back
  • 40348 05/01/00 08:27 Olivier beam gone
  • 40346 05/01/00 08:01 R. Michaels owl shift summary
  • 40344 05/01/00 07:21 R. Michaels beam back
  • 40342 05/01/00 06:59 R. Michaels beam gone
  • 40336 05/01/00 03:25 B. Hu magnet screen
  • 40335 05/01/00 03:24 B. Hu FPP gas strip
  • 40331 05/01/00 01:29 B. Hu FPP eff.